Consider the vectors

Consider the vectors

0 0
Answer #1


3i + 4j - v=-i + 2j-k w=31-31 + 2k (a) Required projection-Z_(+2j-k) (3i+4 (3i + 4j) 5(3i+ 4j 3., 4 916 5 5 b) u # λw and it-